Back to 3.3 Two Basic Rules of Probability

3.3 Exercise 21

Avg difficulty 2.5

Refer to the information in the example. P = tests positive.

  1. Given that a woman develops the disease, what is the probability that she tests positive? Find P(P|B) = 1 − P(N|B).
  2. What is the probability that a woman develops the disease and tests positive? Find P(B AND P) = P(P|B)P(B).
  3. What is the probability that a woman does not develop the disease? Find P(B′) = 1 − P(B).
  4. What is the probability that a woman tests positive for the disease? Find P(P) = 1 − P(N).

OpenStax Statistics, 3.3 Two Basic Rules of Probability, exercise 21

AI Tutor

Adapted from Statistics © 2022 Rice University. Licensed under CC BY 4.0. Source: Statistics. Imported at revision 7dea80a.